OpenStudy (anonymous):

SOLVE. u2 + 2u − 4 = 0

OpenStudy (jamesj):

well, do you know the quadratic formula?

OpenStudy (anonymous):

Do you mean:\[u^2 + 2u - 4 = 0\]

OpenStudy (anonymous):

Yes I meant u^2+2u-4=0. I am lost with this one!

OpenStudy (jamesj):

do you know that quadratic formula?

OpenStudy (anonymous):

You can solve this with the quadratic formula. \[(-b \pm \sqrt{b^2 - 4(a)(c)}) / 2(a)\]

OpenStudy (migitmack):

u=−2+/−22−4(1)(−4)−−−−−−−−−−−√/(2) =−2+/−20−−√/(2) =−2+/−4×5−−−−√/(2) =−2+/−25√/(2) factor out a 2 from the top =2(−1+/−5√)/(2) divide the top and bottom by 2 u=−1+/−5√

OpenStudy (anonymous):

No James, I do not. He didn't give it to us and his example confuses me!

OpenStudy (anonymous):

-2 +/- sqrt(2^2 - 4(1)(-4)) (-2 +/- sqrt(20))2 = (-2 +/- sqrt(4 x 5)) / 2 (-2 +/- 2(sqrt)(5)) / 2 -1 +/- 1(sqrt)(5) \[-1 \pm \sqrt[1]{5}\]

OpenStudy (anonymous):

That's your answer! :)